OpenStudy (anonymous):

Fill in the blank with the numerator of the difference shown below. x+2/x-3 - x+1/x-3

OpenStudy (anonymous):

??/x-3

OpenStudy (pfenn1):

\[\frac{x+2}{x-3} - \frac{x+1}{x-3}=\frac{x+2-x-1}{x-3}\]Can you finish?

OpenStudy (anonymous):

i dont see what you wrote.

OpenStudy (pfenn1):

I wrote (x+2)/(x-3) - (x+1)/(x-3) = [(x+2)-(x+1)]/(x-3)

OpenStudy (anonymous):

should i solve the numerator or keep as is

OpenStudy (anonymous):

*simplify

OpenStudy (pfenn1):

No, simplify, definitely.

OpenStudy (anonymous):

x^+1 ?

OpenStudy (anonymous):

im sorry im just terrible at this lol

OpenStudy (pfenn1):

I'm not sure what that is? (x+2)-(x+1)=x+2-x-1=?

OpenStudy (pfenn1):

I'm only showing the numerator in this last comment. The answer will be over x-3.

OpenStudy (anonymous):

why is 1 added to 2 in the numberator when it says to subtract

OpenStudy (pfenn1):

I'm not sure what you are asking?

OpenStudy (anonymous):

thankyou @pfenn1 :)

OpenStudy (pfenn1):

You're welcome.

OpenStudy (anonymous):

its 1